OpenStudy (anonymous):

What is the solution set for |2x – 8| greater than or equal to 10?

jhonyy9 (jhonyy9):

I2x-8I>= 10 2x-8>=10 2x>=18 x>=18/2 x>=9 and 2x-8<=-10 2x<=-2 x<=-1 x <=-1 x >=9